I can't get an anwser for this
Given
f(x) = f(x) = square root of x - 3
find f^-1(x)

Answer by venugopalramana(3286)

I can't get an anwser for this
Given
f(x) = f(x) =Y= square root of x - 3
find f^-1(x)
X^0.5=Y+3
X=(Y+3)^2
F^-1(X)={F(X)+3}^2